7 of 13 > Determine the (OH), pH, and pOH of a solution with a H+] of 0.011 M at 25 °C. (OH) = pOH =
Determine the H*), pH, and pOH of a solution with an (OH) of 2.7 x 10 Mat 25 °C. pH = pOH =
Determine the H . OH, and pOH of a solution with a pH of 4.40 at 25 °C. [OH-] = pOH =
Determine the H*), OH)and pH of a solution with a pOH of 2.63 at 25 °C. H+) = (OH) - pH =
Answer #1

1)

use:

[OH-] = Kw/[H+]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(1.1*10^-2)

[OH-] = 9.091*10^-13 M

use:

pH = -log [H+]

= -log (1.1*10^-2)

= 1.9586

use:

pOH = -log [OH-]

= -log (9.091*10^-13)

= 12.0414

Answers:

[OH-] = 9.1*10^-13

pH = 1.96

pOH = 12.04

2)

use:

[H+] = Kw/[OH-]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[H+] = (1.0*10^-14)/[OH-]

[H+] = (1.0*10^-14)/2.7*10^-6

[H+] = 3.704*10^-9 M

use:

pH = -log [H+]

= -log (3.704*10^-9)

= 8.4314

use:

pOH = -log [OH-]

= -log (2.7*10^-6)

= 5.5686

Answers:

[H+] = 3.7*10^-9

pH = 8.43

pOH = 5.57

3)

POH = 14 - pH

= 14 - 4.4

= 9.6

use:

pH = -log [H+]

4.4 = -log [H+]

[H+] = 3.981*10^-5 M

use:

pOH = -log [OH-]

9.6 = -log [OH-]

[OH-] = 2.512*10^-10 M

Answers:

[H+] = 4.0*10^-5

[OH-] = 2.5*10^-10

pOH = 9.60

4)

use:

PH = 14 - pOH

= 14 - 2.63

= 11.37

use:

pH = -log [H+]

11.37 = -log [H+]

[H+] = 4.266*10^-12 M

use:

pOH = -log [OH-]

2.63 = -log [OH-]

[OH-] = 2.344*10^-3 M

Answers:

[H+] = 4.3*10^-12

[OH-] = 2.3*10^-3

pH = 11.37